OpenStudy (anonymous):

It takes 40 oz of grass seed to seed 2400 ft^2 of lawn. At this rate, how much would be needed for 6600 ft^2 of lawn?

OpenStudy (andrewhaze):

are you new @shelbynicole09

OpenStudy (anonymous):

I think 110 oz

OpenStudy (anonymous):

because ever 60 ft its one ounce

OpenStudy (legends):

First we have to know how many times does 2400 go into 6600

OpenStudy (andrewhaze):

yup i got the same as @Twigs97

OpenStudy (anonymous):

2400*3 =7200 7200 - 6600= 600 600/ 60 = 10 120 - 10 = 110

OpenStudy (legends):

Lol do you get it @shelbynicole09

OpenStudy (andrewhaze):

yes give @Twigs97 the medal

OpenStudy (anonymous):

yes @Twigs97 is right